Please help to solve :
If one gallon of stain covers 350 square feet, what is the least number of gallons of stain you need to cover the floor of a rectangular deck that is 25 feet long and 16 feet wide?

Answer by checkley75(3666)

area=25*16=400 square feet to paint.
1 galon covers 350 thus 400 needs
400/350=1.14 gallons of paint.